# Data-Driven Discovery of Digital Twins in Biomedical Research

### v1

@@ 2.2.2 Sparse regression
Figure 3 provides an overview of the reviewed sparse regression approaches, which we propose to group into five main families. First, classical sparse regression methods, such as SINDy, infer governing equations from derivative approximations, using libraries of nonlinear candidate terms. This strategy has inspired many extensions aiming to improve SINDy robustness, interpretability, or model selection. For instance, SINDy-AIC (Mangan et al., 2017) outputs multiple models and selects the best ones using the Akaike Information Criterion to balance complexity and goodness-of-fit. It was shown to reliably recover true models for systems like SEIR epidemics (3 variables), where SINDy-AIC identified the correct 6-term model. Next, sparse reaction network recovery includes methods that are specifically amenable for the recovery of biochemical or gene regulatory network structures from time series data. A representative method is Reactive SINDy (Hoffmann et al., 2019), which constructs a library of candidate reactions following mass-action kinetics, enforcing chemical mass conservation that demonstrated effectiveness in inferring gene regulation networks in E. coli and of the MAPK pathway (9 variables). Reactive SINDy successfully recovered the full reaction scheme, retrieving all of the true reactions. Next, sparse regression methods without derivatives approximation has been developed to bypass the sensitivity of numerical differentiation to data noise and sparse sampling. These methods reformulate the regression problem to avoid direct estimation of derivatives, using numerical integration or weak formulations. Next, Bayesian sparse regression enables the use of sparsity-enforcing parameter priors such as spike-and-slab or horseshoe (Piironen and Vehtari, 2017), and help quantify uncertainty of inferred mathematical terms. For example, Jiang et al. (2022) apply a regularized horseshoe prior and MCMC sampling, and successfully recovered synthetic biological networks. Finally, neural sparse regression techniques are increasingly used as they offer to leverage neural networks or LLMs for either estimating derivatives, modeling hidden components, or encoding prior knowledge to guide model search.
@@
Despite many algorithmic variants of the popular SINDy framework, systematic benchmarking was lacking until Kaptanoglu et al. (2023) introduced the first large-scale study using 70 three- or four-dimensional polynomial models of the standardized DYSTS database of chaotic dynamical systems. Their framework enables comparisons of five optimization strategies (i.e., STLSQ (Brunton et al., 2016a), Lasso (Tibshirani, 1996), SR3 (Zheng et al., 2018), MIOSR (Bertsimas and Gurnee, 2023), weak SINDy (Messenger and Bortz, 2021)), under varying data noise levels. Notably, their results emphasize the strong performance of STLSQ and MIOSR and highlight the benefits of the weak SINDy formulation, both in terms of accuracy (low ERMSE) and stability (low variance of ERMSE across noisy experiments). However, this benchmark focuses solely on low-dimensional polynomial systems with full-state observability, which may not capture real-world challenges. Similarly, most studies introducing new model learning methods only compare their approach to one or two earlier baselines (typically STLSQ, Lasso) with limited hyperparameter tuning or statistical evaluation of the performances, reducing the reliability of comparisons. Critically, no benchmark yet evaluates sparse regression methods of all five categories or uses biologically realistic high-dimensional systems documented through irregular data.
@@ 3.1 Handling Noisy, Sparse, and Irregularly Sampled Biological Time Series
A first family of methods aims to reconstruct smooth, continuous trajectories from incomplete or noisy data. This preprocessing step enables the estimation of missing values with simple techniques such as spline-based interpolation and smoothing, low-pass or Savitzky-Golay filters, or kernel regression (Wei, 2022; Bansal et al., 2006; Sun et al., 2022, 2023; Qian et al., 2022; Long et al., 2024; Delahunt and Kutz, 2022; Lejarza and Baldea, 2022). More advanced methods rely on Gaussian processes, which combine smoothing, interpolation, and uncertainty quantification (Atkinson et al., 2019; Bhouri and Perdikaris, 2022; Course and Nair, 2023; Cranmer, 2023; Meng and Qiu, 2025), or on neural networks trained to denoise or interpolate sparse observations (Nardini et al., 2020; Qiu et al., 2024; Yu et al., 2025a; Wu et al., 2025). Other strategies leverage an assumed or partially known model to reconstruct plausible trajectories. For instance, local Taylor approximations generate intermediate states by expanding around available data points ((He et al., 2022)) whereas trajectories may be generated from first-principle models (Abdullah et al. (2022a)). These denoising and reconstruction procedures may be crucial for improving the performance of model inference but may be manipulated with care to avoid erasing biologically-relevant information from data or creating artifacts.
@@
A second category addresses data inconsistencies directly during model inference by adjusting the optimization procedure. For example, Champion et al. (2020) proposed a trimming functionality that automatically down-weights or removes data points with large residuals—often outliers or cases with missing replicates—allowing the algorithm to focus on reliable data. Such approach immediately raises the question of correctly tuning hyperparameters to avoid losing biologically-sound data points and associated information. Following another rationale, Omejc et al. (2024) computes model error only over observed components, ignoring unmeasured data points individually in the loss functions. North et al. (2022) further generalizes this by introducing a time-dependent observation mask that indicates which variables are measured at each time point, enabling flexible integration of partial trajectories. Other approaches directly address data uncertainty in the learning process such as the studies by Hokanson et al. (2023) and Kaheman et al. (2022) that propose to jointly estimate a clean version of the data and the dynamic model.
@@
Biological challenge 1 - Key takeaways Overall, handling incomplete or noisy biological data may either be performed through an initial step of data reconstruction or denoising, or by a reformulation of the optimization procedure for model inference. Data smoothing and imputation are often critical for ensuring that inferred models reflect the underlying dynamics, especially in the presence of limited or corrupted observations. However, such procedures must be well adapted to the biological problem and raise the issue of hyperparameter tuning to avoid losing information or incorrectly transforming data.
@@ 3.2 Leveraging Data in Multiple Experimental Conditions or Individuals
Several methods leverage neural networks to discover or generalize shared dynamical structures. NeuralCODE (Huang et al., 2025b) uses Automated Machine Learning (AutoML), i.e., automated architecture search, to design neural network architectures that separate shared from subject-specific parameters, effectively learning a common backbone of differential equations for biological processes consistent across patients, with personalized adjustments. In the same direction, Gui et al. (2025) propose moving beyond explaining variability through parameter shifts alone via invariant function learning, allowing both parameters and functional forms to vary across conditions. A neural network estimates the invariant component of the ODE, enabling, as a second step, symbolic or sparse regression to recover interpretable models for each subject. MetaPhysiCa (Mouli et al., 2023) casts dynamics learning as a meta-learning problem, identifying physical laws that generalize across experimental setups while adapting quickly to new conditions, like unseen initial states or parameters. By combining causal inference and invariant risk minimization, it extracts robust, generalizable models even in extrapolation regimes. Complementing these approaches, OASIS (Bhadriraju et al., 2020) pairs SINDy with a neural network that maps input conditions to ODE parameters. Local models are learned offline from condition-specific time series, while the network generalizes across them, enabling rapid adaptation to new regimes via online parameter prediction. Given their reliance on neural networks, these methods likely require larger amounts of data per patient to ensure accurate learning and generalization.
@@ 3.3 Integrating Prior Biological Knowledge into Model Discovery
Prior knowledge may be defined as any supplementary information that helps constrain the otherwise vast and ill-posed problem of inferring ODEs from biological time series—such as known species interactions, kinetic forms, or plausible parameter values.
@@ 3.3.1 Prior knowledge on chemical interactions
Prior knowledge about species interactions explicitly involves information regarding the possibility of a reactant $R$ becoming a product $P$. Such knowledge can be incorporated at the whole model level or reaction level, as described below.
Model-level priors encode global structural assumptions, such as sparsity —favoring minimal interaction sets— or topological features like scale-free networks where most nodes have few connections and a few act as hubs (Barabási and Albert, 1999; Ouma et al., 2018). These assumptions are commonly enforced in sparse regression frameworks (see Section 2.2.2). For instance, SINDy (Brunton et al., 2016a) uses $\ell_{1}$ regularization (LASSO Tibshirani (1996)), while SR3 (Champion et al., 2020) applies the sparsity penalty to an auxiliary variable rather than directly to the coefficients, improving optimization stability. Bayesian approaches allow imposing flexible priors on parameters, such as the regularized horseshoe, which shrinks most coefficients to zero while allowing a few large ones, or the spike-and-slab, which explicitly models inclusion via a mixture of near-zero (’spike’) and unconstrained (’slab’) components (Jiang et al., 2022; Bhouri and Perdikaris, 2022; Nayek et al., 2021). While accounting for such priors provides convex relaxations of the nonconvex model selection problem, Mixed-Integer Optimization (MIO) offers a more principled alternative by explicitly controlling term inclusion via binary indicators that encode whether a given term is included (1) or excluded (0) from the model. As an important example of such method, MIOSR (Bertsimas and Gurnee, 2023) formulates symbolic model discovery as a constrained optimization problem, enabling provably optimal recovery of sparse ODE systems from noisy data. Alternatively, Reactmine (Martinelli et al., 2022) tackles model complexity through a tree-search algorithm that infers reactions sequentially, using a depth limit to enforce parsimony by assuming few reactions are sufficient to explain the data.
@@ 3.3.2 Prior knowledge on mathematical formulation or parameters of interaction kinetics
Kinetic prior knowledge refers to assumptions or known forms of reaction kinetics, i.e., mathematical expressions defining reaction rates, and associated parameter values. This can apply to the whole model or to specific interactions.
@@ 3.3.3 Foundation models-aided integration of prior knowledge
Beyond direct inclusion of prior knowledge, the advent of Foundation Models (FMs) and LLMs offer new opportunities to incorporate more abstract forms of available information into model inference. Pre-trained on large-scale, heterogeneous corpora like scientific texts, databases, and web contents, these models internalize a wide range of domain knowledge, recognizing patterns and conventions that would otherwise require manual encoding. In biology, for instance, specialized FMs like scGPT (Cui et al., 2024) and scFoundation (Hao et al., 2024) can generate single-cell transcriptomics data, as a result of their training on millions of gene expression profiles collected across diverse tissues and conditions, which allowed them to implicitly capture biological structures such as cell-type hierarchies and gene co-expression. Importantly, such tools may assist tasks like model initialization or constraint suggestion.
@@
Biological challenge 3 - Key takeaways Despite rapid progress, the integration of prior knowledge into biological model discovery remains underdeveloped and fragmented. While methods incorporating topological constraints and known mathematical terms or parameter priors into digital twin inference have recently been developed, this is often done in an ad hoc or inflexible manner, and the automatic integration of structured large-scale biological databases remain challenging due to mapping difficulties and context sensitivity. Foundation Models and LLMs offer an intriguing new avenue for bridging these gaps, yet with limitations mainly due to their current lack of reliability that would deserve more in-depth evaluation.
@@ 3.4 Dealing with Latent or Unobserved Variables
Another approach developed by Somacal et al. (2022) proposes to regress higher-order derivatives of observed variables onto a dictionary of candidate functions, allowing the influence of hidden components to be encoded through derivative interactions. This leverages the fact that, under certain assumptions, unobserved states leave signatures in the derivatives of observed variables (Takens, 2006). In the same line, Martinelli et al. (2021) models the indirect action of a variable $z$ through its time-integrated effect, $Z(t)=k\int_{0}^{t}z(s)\mathrm{d}s$, which assumes that its influence is mediated through an intermediate species, provided that all reactions follow the law of mass action. Interestingly, this enables modeling delayed - rather than instantaneous - regulatory effects when only upstream regulators are documented. Overall, indirect strategies like encoding latent influence via higher-order derivatives or integral effects offer elegant workarounds but typically rely on strong assumptions holding in a limited number of cases.
A last class of approaches focuses on reconstructing latent states using neural networks. Lu et al. (2022a) propose a framework where a neural encoder maps temporal sequences of observed variables to estimates of the unobserved components. From these reconstructed states, governing equations are then identified by fitting a sparse combination of predefined candidate functions. Grigorian et al. (2025) similarly train a hybrid neural ODE model on the observed trajectories and apply symbolic regression to the latent part of the learned representation, yielding interpretable equations. These approaches combine the expressive power of neural networks with the interpretability of symbolic models, enabling recovery of hidden dynamics from partial observations. Yet, major challenges remain. Methods that explicitly infer latent trajectories often suffer from identifiability issues, making inference sensitive to initialization and priors. Neural reconstruction techniques, while flexible, may obscure mechanistic interpretability and can overfit in low-data regimes.
@@ 4.1 High Dimension of Biological Systems leading to computational and low-data regime challenge
These challenges are twofold. First, discovering larger systems requires high computational resources: as model dimension increases, sparse regression algorithms must handle rapidly growing candidate function libraries, while symbolic methods face a combinatorial explosion of possible expressions. For instance, performing sparse regression for a $10$-variable system and a function library $\bm{\Theta}(\mathbf{Y})$ containing first- and second-order terms yields 65 possible terms; selecting up to 4 active terms per equation already produces over 722,000 possible combinations per equation. Second, high dimensionality requires more numerous and diverse datasets to constrain the model and avoid overfitting, which may be challenging considering the cost of biological experiments. However, with the rise of high-throughput technologies such as bulk and single-cell multi-omics profiling, genetic and drug screening, or perturbation platforms like CRISPR-based perturb-seq, collecting large-scale time-varying datasets is becoming increasingly feasible.
@@
Methodological challenge 1 - Key takeaways While promising strategies exist, most methods still struggle beyond modest system sizes. A key challenge is to design scalable algorithms that retain interpretability and mechanistic accuracy in high dimensions. This likely requires hybrid approaches that combine structural priors, modularity, and efficient inference to navigate vast model spaces without exhaustive search. Equally crucial is leveraging existing databases for incorporating available prior knowledge and systematically incorporating large-scale datasets like multi-omics profiling or genetic screens, which provide rich perturbation data for consistent dynamics recovery (Section 3.2).

### v2

@@ 2.2.2 Sparse regression
Figure 3 provides an overview of the reviewed sparse regression approaches, which we propose to group into five main families. First, classical sparse regression methods, such as SINDy, infer governing equations from derivative approximations, using libraries of nonlinear candidate terms. This strategy has inspired many extensions aiming to improve SINDy robustness, interpretability, or model selection. For instance, SINDy-AIC (Mangan et al., 2017) outputs multiple models and selects the best ones using the Akaike Information Criterion to balance complexity and goodness-of-fit. It was shown to reliably recover true models for systems like SEIR epidemics (3 variables), for which itidentified the correct 6-term model. Next, sparse reaction network recovery includes methods that are specifically amenable for biochemical or gene regulatory network structures from time series data. A representative method is Reactive SINDy (Hoffmann et al., 2019), which constructs a library of candidate reactions following mass-action kinetics, enforcing chemical mass conservation that retrieved all true reactions of the E. coli and of MAPK pathway networks (respectively, 3 and 9 variables). Alternatively, sparse regression methods without derivatives approximation has been developed to bypass the sensitivity of numerical differentiation to data noise and sparse sampling. These methods reformulate the regression problem to avoid direct estimation of derivatives, using numerical integration or weak formulations. Next, Bayesian sparse regression enables the use of sparsity-enforcing parameter priors such as spike-and-slab or horseshoe (Piironen and Vehtari, 2017), and help quantify uncertainty of inferred mathematical terms. For example, Jiang et al. (2022) apply a regularized horseshoe prior and MCMC sampling, and successfully recovered synthetic biological networks. Finally, neural sparse regression techniques are increasingly used as they offer to leverage neural networks or LLMs for either estimating derivatives, modeling hidden components, or encoding prior knowledge to guide model search.
@@
Despite many algorithmic variants of the popular SINDy framework, systematic benchmarking was lacking until Kaptanoglu et al. (2023) introduced the first large-scale study using 70 three- or four-dimensional polynomial models of the standardized DYSTS database of chaotic dynamical systems. Their framework enables comparisons of five optimization strategies (i.e., STLSQ (Brunton et al., 2016a), Lasso (Tibshirani, 1996), SR3 (Zheng et al., 2018), MIOSR (Bertsimas and Gurnee, 2023), weak SINDy (Messenger and Bortz, 2021)), under varying data noise levels. Notably, their results emphasize the strong performance of STLSQ and MIOSR and highlight the benefits of the weak SINDy formulation, both in terms of accuracy (low RMSE) and stability (low variance of RMSE across noisy experiments). However, this benchmark focuses solely on low-dimensional polynomial systems with full-state observability, which may not capture real-world challenges. Similarly, most studies introducing new model learning methods only compare their approach to one or two earlier baselines (typically STLSQ, Lasso) with limited hyperparameter tuning or statistical evaluation of the performances, reducing the reliability of comparisons. Critically, no benchmark yet evaluates sparse regression methods of all five categories or uses biologically realistic high-dimensional systems documented through irregular data.
@@ 3.1 Handling Noisy, Sparse, and Irregularly Sampled Biological Time Series
A first family of methods aims to reconstruct smooth, continuous trajectories from incomplete or noisy data. This preprocessing step enables the estimation of missing values with simple techniques such as spline-based interpolation and smoothing, low-pass or Savitzky-Golay filters, or kernel regression (Wei, 2022; Bansal et al., 2006; Sun et al., 2022, 2023; Qian et al., 2022; Long et al., 2024; Delahunt and Kutz, 2022; Lejarza and Baldea, 2022). More advanced methods rely on Gaussian processes, which combine smoothing, interpolation, and uncertainty quantification (Atkinson et al., 2019; Bhouri and Perdikaris, 2022; Course and Nair, 2023; Cranmer, 2023; Meng and Qiu, 2025), or on neural networks trained to denoise or interpolate sparse observations (Nardini et al., 2020; Qiu et al., 2024; Yu et al., 2025a; Wu et al., 2025). Other strategies leverage an assumed or partially known model to reconstruct plausible trajectories. For instance, local Taylor approximations generate intermediate states by expanding around available data points ((He et al., 2022)) whereas trajectories may be generated from first-principle models (Abdullah et al. (2022a)). These denoising and reconstruction procedures may be crucial for improving the performance of model inference but must be manipulated with care to avoid erasing biologically-relevant information from data or creating artifacts.
@@
A second category addresses data inconsistencies directly during model inference by adjusting the optimization procedure. For example, Champion et al. (2020) proposed a trimming functionality that automatically down-weights or removes data points with large residuals—often outliers or cases with missing replicates—allowing the algorithm to focus on reliable data. Such approach immediately raises the question of correctly tuning hyperparameters to avoid losing biologically-sound data points and associated information. Following another rationale, Omejc et al. (2024) restrict the computation of the model error to observed components, at documented time points. North et al. (2022) further generalizes this by introducing a time-dependent observation mask that indicates which variables are measured at each time point, enabling flexible integration of partial trajectories. Other approaches directly address data uncertainty in the learning process such as the studies by Hokanson et al. (2023) and Kaheman et al. (2022) that propose to jointly estimate a clean version of the data and the dynamic model.
@@
Biological challenge 1 - Key takeaways Overall, handling incomplete or noisy biological data may either be performed through an initial step of data reconstruction or denoising, or by an adaptation of the optimization procedure for model inference. Data smoothing and imputation are often critical for ensuring that inferred models reflect the underlying dynamics, especially in the presence of limited or corrupted observations. However, such procedures must be well adapted to the biological problem and raise the issue of hyperparameter tuning to avoid losing information or incorrectly transforming data.
@@ 3.2 Leveraging Data in Multiple Experimental Conditions or Individuals
Several frameworks leverage neural networks to discover or generalize shared dynamical structures. NeuralCODE (Huang et al., 2025b) uses Automated Machine Learning (AutoML), i.e., automated architecture search, to design neural network architectures that separate shared from subject-specific parameters, effectively learning a common backbone of differential equations for biological processes consistent across patients, with personalized adjustments. In the same direction, Gui et al. (2025) propose moving beyond explaining variability through parameter shifts alone via invariant function learning, allowing both parameters and functional forms to vary across conditions. A neural network estimates the invariant component of the ODE, enabling, as a second step, symbolic or sparse regression to recover interpretable models for each subject. Alternatively, MetaPhysiCa (Mouli et al., 2023) casts dynamics learning as a meta-learning problem, identifying physical laws that generalize across experimental setups while efficiently adapting to new conditions, like unseen initial states or parameters. By combining causal inference and invariant risk minimization, it extracts robust, generalizable models even in extrapolation regimes. Complementing these approaches, OASIS (Bhadriraju et al., 2020) pairs SINDy with a neural network that maps input conditions to ODE parameters. Local models are learned offline from condition-specific time series, while the network generalizes across them, enabling rapid adaptation to new regimes via online parameter prediction. Given their reliance on neural networks, these methods likely require larger amounts of data per patient to ensure accurate learning and generalization.
@@ 3.3 Integrating Prior Biological Knowledge into Model Discovery
Prior knowledge may be defined as any supplementary information that helps constrain the otherwise vast and ill-posed problem of inferring equations governing a biological system from time series—such as known species interactions, kinetic mathematical forms, or plausible parameter values.
@@ 3.3.1 Prior knowledge on chemical interactions
Prior knowledge about species interactions explicitly involves information regarding the possibility of a reactant $R$ becoming a product $P$. Such knowledge can be incorporated at the whole model level or at the reaction level, as described below.
Model-level priors encode global structural assumptions, such as sparsity —favoring minimal interaction sets— or topological features like scale-free networks where most nodes have few connections and a few act as hubs (Barabási and Albert, 1999; Ouma et al., 2018). These assumptions are commonly enforced in sparse regression frameworks (see Section 2.2.2). For instance, SINDy (Brunton et al., 2016a) uses $\ell_{1}$ regularization (LASSO, Tibshirani (1996)), while SR3 (Champion et al., 2020) applies the sparsity penalty to an auxiliary variable rather than directly to the coefficients, improving optimization stability. Bayesian approaches allow imposing flexible priors on parameters, such as the regularized horseshoe, which shrinks most coefficients to zero while allowing a few large ones, or the spike-and-slab, which explicitly models inclusion via a mixture of near-zero (’spike’) and unconstrained (’slab’) components (Jiang et al., 2022; Bhouri and Perdikaris, 2022; Nayek et al., 2021). While accounting for such priors provides convex relaxations of the nonconvex model selection problem, Mixed-Integer Optimization (MIO) offers a more principled alternative by explicitly controlling term inclusion via binary indicators that encode whether a given term is included (1) or excluded (0) from the model. As an important example of such method, MIOSR (Bertsimas and Gurnee, 2023) formulates symbolic model discovery as a constrained optimization problem, enabling provably optimal recovery of ODE systems from noisy data. Alternatively, Reactmine (Martinelli et al., 2022) tackles model complexity through a tree-search algorithm that infers reactions sequentially, using a depth limit to enforce parsimony by assuming few reactions are sufficient to explain the data.
@@ 3.3.2 Prior knowledge on mathematical formulation or parameters of interaction kinetics
Kinetic prior knowledge refers to assumptions or known forms of reaction kinetics, i.e., mathematical expressions of the reaction rates and associated parameter values. This can apply to the whole model or to specific interactions.
@@ 3.3.3 Foundation models-aided integration of prior knowledge
Beyond direct inclusion of prior knowledge, the advent of Foundation Models (FMs) and LLMs offer new opportunities to incorporate more abstract forms of available information into model inference. Pre-trained on large-scale, heterogeneous corpora like scientific texts, databases, and web contents, these models internalize a wide range of domain knowledge, recognizing patterns and conventions that would otherwise require manual encoding. In biology, for instance, specialized FMs like scGPT (Cui et al., 2024) and scFoundation (Hao et al., 2024) can generate single-cell transcriptomics data, as a result of their training on millions of gene expression profiles collected across diverse tissues and conditions, which allowed them to implicitly capture biological structures such as cell-type hierarchies and gene co-expression. Importantly for this review, such tools may assist tasks like model initialization or constraint suggestion.
@@
Biological challenge 3 - Key takeaways Despite rapid progress, the integration of prior knowledge into biological model discovery remains underdeveloped and fragmented. While methods incorporating topological constraints and known mathematical terms or parameter priors into digital twin inference have recently been developed, this is often done in an ad hoc or inflexible manner, and the automatic integration of large-scale structured biological databases remain challenging due to mapping difficulties and context sensitivity. Foundation Models and LLMs offer an intriguing new avenue for bridging these gaps, yet with limitations mainly due to their current lack of reliability that would deserve more in-depth evaluation.
@@ 3.4 Dealing with Latent or Unobserved Variables
Another approach developed by Somacal et al. (2022) proposes to regress higher-order derivatives of observed variables onto a dictionary of candidate functions, allowing the influence of hidden components to be encoded through derivative interactions. This leverages the fact that, under certain assumptions, unobserved states leave signatures in the derivatives of observed variables (Takens, 2006). In the same line, Martinelli et al. (2021) models the indirect action of a variable $z$ through its time-integrated effect, $Z(t)=k\int_{0}^{t}z(s)\mathrm{d}s$, which assumes that its influence is mediated through an intermediate species, provided that the reaction follows the law of mass action. Interestingly, this enables modeling delayed - rather than instantaneous - regulatory effects when only upstream regulators are documented. Overall, indirect strategies like encoding latent influence via higher-order derivatives or integral effects offer elegant workarounds but typically rely on strong assumptions holding in a limited number of cases.
A last class of approaches focuses on reconstructing latent states using neural networks. Lu et al. (2022a) propose a framework where a neural encoder maps temporal sequences of observed variables to estimate the unobserved components. From these reconstructed states, governing equations are then identified by fitting a sparse combination of predefined candidate functions. Grigorian et al. (2025) similarly train a hybrid neural ODE model on the observed trajectories and apply symbolic regression to the latent part of the learned representation, yielding interpretable equations. These approaches combine the expressive power of neural networks with the interpretability of symbolic models, enabling recovery of hidden dynamics from partial observations. Yet, major challenges remain. Methods that explicitly infer latent trajectories often suffer from identifiability issues, making inference sensitive to initialization and priors. Neural reconstruction techniques, while flexible, may obscure mechanistic interpretability and can overfit in low-data regimes.
@@ 4.1 High Dimension of Biological Systems leading to computational and low-data regime challenge
These challenges are twofold. First, discovering larger systems requires high computational resources: as model dimension increases, sparse regression algorithms must handle rapidly growing candidate function libraries, while symbolic methods face a combinatorial explosion of possible expressions. For instance, performing sparse regression for a $10$-variable system and a function library $\bm{\Theta}(\mathbf{Y})$ containing first- and second-order terms yields 65 possible terms; selecting up to 4 active terms per equation already produces over 722,000 possible combinations per equation. Second, high dimensionality requires more numerous and diverse datasets to constrain the model and avoid overfitting, which may be challenging considering the cost of biological experiments. However, with the rise of high-throughput technologies such as bulk and single-cell multi-omics profiling, CRISPR-based genetic or drug screening, collecting large-scale time-varying datasets is becoming increasingly feasible.
@@
Methodological challenge 1 - Key takeaways While promising strategies exist, most methods still struggle beyond modest system sizes. A key challenge is to design scalable algorithms that retain interpretability and mechanistic accuracy in high dimension. This likely requires hybrid approaches that combine structural priors, modularity, and efficient inference to navigate vast model spaces without exhaustive search. Equally crucial is leveraging existing databases for systematically incorporating large-scale datasets like multi-omics profiling or genetic screens , which provide rich perturbation data constraining dynamics recovery (Section 3.2).